A train averages a speed of 90 miles per hour across the plains and 37.5 miles per hour through the mountains. If a trip of 300 miles took 3 hour and 48 minutes, how much of it was through the mountains?
Answer by fractalier(6550)

Call the time through the mountains, t.
Since 3 hours, 48 minutes is 3 4/5 (or 3.8) hours, the time on the plains must be 3.8 - t.
We can set this up, for rt = d, as
90(3.8 - t) + 37.5t = 300
342 - 90t + 37.5t = 300
-52.5t = -42
t = .8 hours = 48 minutes